# Mechanistic Permutability: Match Features Across Layers

### v2

@@ 2 Background
Here, $\sigma$ is an activation function (e.g., ReLU). Importantly, ${\bm{f}}({\bm{x}})$ is made to be sparse to help the SAE capture monosemantic features. The final loss is defined as $\mathcal{L}({\bm{x}})=\|{\bm{x}}-\hat{{\bm{x}}}({\bm{f}}({\bm{x}}))\|_{2}+\lambda|\|{\bm{f}}({\bm{x}})\|_{0}$. The sparsity of the model can be increased by adjusting the $\lambda$ hyperparameter, which balances the trade-off between the reconstruction term and the regularization term. However, determining the optimal value for $\lambda$ is not straightforward.
@@ 3 Matching Sparse Autoencoder Features
${\bm{P}}^{(A\rightarrow B)}=\underset{{\bm{P}}\in\mathcal{P}_{F}}{\arg\min}\sum_{i=1}^{d}\|{\bm{W}}_{\mathrm{dec}_{i,:}}^{(A)}-{\bm{P}}{\bm{W}}_{\mathrm{dec}_{i,:}}^{(B)}\|_{2}=\underset{{\bm{P}}\in\mathcal{P}_{F}}{\arg\max}\left\langle{\bm{P}},\left({\bm{W}}_{\mathrm{dec}}^{(A)}\right)^{\top}{\bm{W}}_{\mathrm{dec}}^{(B)}\right\rangle_{F},$ (3)
@@ 3.1 Folding Sparse Autoencoder Weights
Our hypothesis is that parameter folding helps align decoder weights in a way that reflects the actual dynamics of hidden state norms during model evaluation. Specifically, we observed that the learned thresholds ${\bm{\theta}}$ capture the growth of hidden state norms across layers (see Figure 1). We observed that ${\bm{\theta}}$ values, in fact, encapsulate the growth of hidden state norms. By folding ${\bm{\theta}}$ into the weights, we normalize the decoder weights to align with these dynamics.
@@
Folding the activation thresholds into the weights aligns decoder columns with the dynamics of hidden state norms. This process allows us to match decoder weights based on their MSE in the space of actual hidden state norms. By folding parameters, we can match features across layers in a data-free manner, without needing input data to capture input scales. For a schematic illustration of hypothetical behavior, refer to Figure 2.
@@ 3.2 Composing Permutations
If layers A and C are not too far apart, we can adequately approximate the matching between them by composing the appropriate permutation matrices from intermediate layers.
@@ 4 Experimental Setup
We conducted experiments using Sparse Autoencoders (SAEs) trained on the Gemma 2 model, with feature descriptions from Neuronpedia generated by GPT-4o. We focused on SAEs with a regularization coefficient around 70 (for a complete list of SAEs used, refer to Appendix Table 1). For matching we used MSE from both decoder and encoder layers. During our initial experiments, we observed that the decoder-only option performs similarly to our scheme, while the encoder-only suffers from poor quality of matching (see Appendix Figure 11 for comparison).
@@
    An external large language model (LLM) compared Neuronpedia descriptions of matched features, categorizing them as "SAME" (identical meanings), "MAYBE" (possibly similar meanings), or "DIFFERENT" (distinct meanings). Each experiment involved approximately 1,600 LLM evaluations over 100 feature paths spanning 16 layers (details in Appendix Section B).
@@ 5.1 Understanding Parameter Folding
As Hypothesis 2 relies on the observation that ${\bm{\theta}}$ incapsulates differences in hidden state norms, we evaluated MSE of obtained permutations. More concretely, we evaluated MSE for folded matching, and for unfolded and naive (without any permutations) matching strategies, for which we first permuted SAE weights based on unfolded MSE values, and then evaluated MSE of appropriate folded weights. See Figure 3 for the results. Initially, naive matching displayed high MSE across all SAE parameters, indicating the lack of assurance in maintaining feature order between layers. Notably, unfolded matching showed higher MSE in the scale of hidden state representations, supporting behaviour described with Hypothesis 2. Also note that since MSE is plotted with folded parameters, the MSE of the Encoder layer decreases, while the MSE of the Decoder layer increases (this behavior is explained by Equation 4).
@@
We compared the performance of folded and unfolded matching based on semantic similarity with an external LLM. The results can be seen in Figure 4 for LLM evaluation across different layers. Folded matching exhibited improved feature matching quality compared to the unfolded variant. Also, we observed a performance drop at the $10$-th layer, which we study in Section 5.3. Combining all the results from this section, we validated Hypothesis 2 regarding the behavior of matching with folded parameters.
@@ 5.2 Features Matching
We also evaluated the distribution of MSE values for three pairs of layers matched with folded parameters, grouped according to external LLM evaluation. The results, shown in Figure 6, similarly revealed that lower MSE values were associated with the semantic similarity of matched features. It is important to note that the distribution of MSE values across different layer pairs varied due to differences in weight scales (as described in Section 3.1).
@@
Figure 7: Left: External LLM evaluation of feature matching across layers with folded matching. Right: Matching Score across layers. The LLM evaluation shows near-zero "SAME" features in the initial layers, increasing after the 10th layer. The Matching Score shows a similar pattern but does not drop to zero in the initial layers, indicating some predictive ability. These results suggest that initial layers may have more polysemantic features, making them harder to evaluate using Neuronpedia descriptions. See Section 5.3 for more details.
@@ 5.3 On the Performance of Matching Across Layers
These observations lead us to hypothesize that the features extracted by the SAEs in the initial layers may suffer from polysemanticity—they activate in response to multiple, unrelated inputs and represent overlapping concepts. The descriptions in Neuronpedia cover only a single mode of these polysemantic features, potentially missing other aspects of their activation patterns. As a result, when the external LLM evaluates semantic similarity based on these limited descriptions, it identifies few "SAME" features in the initial layers.
This phenomenon aligns with findings from previous research. Gurnee et al. (2023) also reported increased polysemanticity in the early layers of neural networks. In the initial layers, features tend to be more general and less specialized, capturing a wide range of inputs. As the information propagates through the network, features become more monosemantic, specializing in specific concepts or functions.
Figure 8: Results of the external LLM evaluation. We compare features from 10th, 12th and 20th layers with matched features in subsequent layers. It can be observed that feature similarity gradually declines, but remains significant for approximately five layers. See Section 5.4 for more details.
@@ 5.5 SAE Match as a Layer Pruning
Figure 9 provides a schematic illustration of this layer pruning method.
We observed the smallest performance drop on the OpenWebText dataset, while a decrease in quality was noted at the last layer of the Code dataset. These findings indicate that the matched features share similar semantics, enabling us to approximate the next hidden states for all layers beyond the 10th layer.
For experiments on how the sparsity level of the SAE affects feature matching, see Appendix Section A.
@@ 6 Limitations and Conclusion
We introduced SAE Match, a novel data-free method for aligning Sparse Autoencoder (SAE) features across neural network layers, tackling the challenge of understanding feature evolution amid polysemanticity and feature superposition. By minimizing the mean squared error between folded SAE parameters—which integrate activation thresholds into the weights—we accounted for feature scale differences and improved matching accuracy. Experiments on the Gemma 2 language model validated our approach, showing effective feature matching across layers and that certain features persist over multiple layers. Additionally, we demonstrated that our method enables approximating hidden states across layers. This advances mechanistic interpretability by providing a new tool for analyzing feature dynamics and deepening our understanding of neural network behavior.
While we evaluated our method using SAEs from the Gemma 2 model with promising results, further research is needed to confirm if these findings generalize to other models. Additionally, potential errors in Neuronpedia feature descriptions may affect the precise evaluation of semantic similarity.
@@ Appendix A On the Matching with Different Sparsity
Appendix B Evaluation Details
@@
0 width_16k/average_l0_46 13 width_16k/average_l0_83
1 width_16k/average_l0_40 14 width_16k/average_l0_83
@@
12 width_16k/average_l0_82 25 width_16k/average_l0_55
Table 1: Full list of SAEs we used for our experiments. To the best of our knowledge these SAEs are used in Neuronpedia website.
Appendix C Path Examples
@@
Figure 13: Examples of paths first, second, third and fourth features from 20th layer to 25th layer.

### v3

@@ 2 Background
Here, $\sigma$ is an activation function (e.g., ReLU). Importantly, ${\bm{f}}({\bm{x}})$ is made to be sparse to help the SAE capture monosemantic features. The final loss is defined as $\mathcal{L}({\bm{x}})=\operatorname{MSE}({\bm{x}},\hat{{\bm{x}}}({\bm{f}}({\bm{x}})))+\lambda\mathcal{L}_{reg}$. The sparsity of the model can be increased by adjusting the $\lambda$ hyperparameter, which balances the trade-off between the reconstruction term and the regularization term. However, determining the optimal value for $\lambda$ is not straightforward.
@@ 3 Matching Sparse Autoencoder Features
${\bm{P}}^{(A\rightarrow B)}=\underset{{\bm{P}}\in\mathcal{P}_{F}}{\arg\min}\sum_{i=1}^{d}\|{\bm{W}}_{\mathrm{dec}_{i,:}}^{(A)}-{\bm{P}}{\bm{W}}_{\mathrm{dec}_{i,:}}^{(B)}\|^{2}=\underset{{\bm{P}}\in\mathcal{P}_{F}}{\arg\max}\left\langle{\bm{P}},\left({\bm{W}}_{\mathrm{dec}}^{(A)}\right)^{\top}{\bm{W}}_{\mathrm{dec}}^{(B)}\right\rangle_{F},$ (3)
@@ 3.1 Folding Sparse Autoencoder Weights
Our hypothesis is that parameter folding helps align decoder weights in a way that reflects the actual dynamics of hidden state norms during model evaluation (see Figure 1). We observed that ${\bm{\theta}}$ values, in fact, encapsulate the growth of hidden state norms. By folding ${\bm{\theta}}$ into the weights, we normalize the decoder weights to align with these dynamics. This process allows us to match decoder weights based on their MSE in the space of actual hidden state norms. By folding parameters, we can match features across layers in a data-free manner, without needing input data to capture input scales. For a schematic illustration of behavior, refer to Figure 2
@@
Folding the activation thresholds into the weights improves the quality of matching.
@@ 3.2 Composing Permutations
We can approximate the matching between any two layers, A and C, by composing permutations. However, the quality of this approximation decreases as the relative distance between these layers increases.
@@ 4 Experimental Setup
We conducted experiments using Sparse Autoencoders (SAEs) trained on the Gemma 2 model, with feature descriptions from Neuronpedia generated by GPT-4o. We focused on SAEs with descriptions provided by Neuronpedia (for a complete list of SAEs used, refer to Appendix Table 1). For matching we used MSE from both decoder and encoder layers. During our initial experiments, we observed that the decoder-only option performs similarly to our scheme, while the encoder-only suffers from poor quality of matching (see Appendix Figure 12 for comparison). A more detailed analysis of the differences in matching using various sets of weights is deferred to future work.
@@
    An external large language model (LLM) compared Neuronpedia descriptions of matched features, categorizing them as "SAME" (identical meanings), "MAYBE" (possibly similar meanings), or "DIFFERENT" (distinct meanings). Each experiment involved approximately 1,600 LLM evaluations over 100 feature paths spanning 16 layers (details in Appendix Section C).
@@ 5.1 Understanding Parameter Folding
As Hypothesis 2 relies on the observation that ${\bm{\theta}}$ incapsulates differences in hidden state norms, we evaluated MSE of obtained permutations. More concretely, we evaluated MSE for folded matching, and for unfolded and naive (without any permutations) matching strategies, for which we first permuted SAE weights based on unfolded MSE values, and then evaluated MSE of appropriate folded weights. See Figure 3 for the results. Initially, naive matching displayed high MSE across all SAE parameters, indicating the lack of assurance in maintaining feature order between layers. Unfolded matching showed higher MSE in the scale of hidden state representations. Also note that since MSE is plotted with folded parameters, the MSE of the Encoder layer decreases, while the MSE of the Decoder layer increases (this behavior is explained by Equation 4).
@@
We compared the performance of folded and unfolded matching based on semantic similarity with an external LLM. The results can be seen in Figure 4 for LLM evaluation across different layers. Folded matching exhibited improved feature matching quality compared to the unfolded variant. This result supports Hypothesis 2 regarding the behavior of matching with folded parameters.
Also, we observed a performance drop at the $10$-th layer, which we study in Section 5.3.
@@ 5.2 Features Matching
We also evaluated the distribution of MSE values for three pairs of layers matched with folded parameters, grouped according to external LLM evaluation. The results, shown in Figure 6, similarly revealed that lower MSE values were associated with the semantic similarity of matched features. It is important to note that the distribution of MSE values across different layer pairs varied due to differences in weight scales (as described in Section 3.1). See Appendix Section B for experiments with MSE thresholds.
@@
Figure 7: Left: External LLM evaluation of feature matching across layers with folded matching. Right: Matching Score across layers and across different matching methods. Folding parameters has a positive effect on Matching Score. The LLM evaluation shows near-zero "SAME" features in the initial layers, increasing after the 10th layer. The Matching Score shows a similar pattern but does not drop to zero in the initial layers, indicating some predictive ability. These results suggest that initial layers may have more polysemantic features, making them harder to evaluate using Neuronpedia descriptions. See Section 5.3 for more details.
@@ 5.3 On the Performance of Matching Across Layers
Based on these results, SAE Match fails to align similar features in the initial layers. Our main explanation for this phenomenon lies in the differences in $l_{0}$norms at these early layers (see Figure 15 for more details). We observe that when utilizing canonical SAEs, modules in the initial layers exhibit large variance in $l_{0}$ values. If we utilize SAEs with similar norms at each layer, the Explained Variance improves significantly. While we continue our experiments with canonical SAEs for consistency, this result suggests that these layers should be selected more precisely in future studies.
Figure 8: Results of the external LLM evaluation. We compare features from 10th, 12th and 20th layers with matched features in subsequent layers. It can be observed that feature similarity gradually declines, but remains significant for approximately five layers. We use 1 for DIFFERENT features, 2 for MAYBE features, and 3 for SAME features, and then average this score across layers. See Section 5.4 for more details.
@@ 5.5 SAE Match as a Layer Pruning
Figure 11: Left: The change in cross-entropy loss ($\Delta L$) with each layer of the SAE. Right: The explained variance of the approximated hidden states for each layer of the SAE. We present results for plain encoding and decoding of the current and previous hidden states, as well as for matching in both folded and unfolded variants. While folded matching eventually matches the results of encoding and decoding the previous hidden state, at lower layers it introduces errors that reduce performance.
Figure 9 provides a schematic illustration of this layer pruning method. See Figure 10 for the results.
We observed the smallest performance drop on the OpenWebText dataset, while a decrease in quality was noted at the last layer of the Code dataset. These findings indicate that the matched features share similar semantics, enabling us to approximate the next hidden states for all layers beyond the 10th layer. Also see Figure 11 for the Explained Variance and the change in loss across layers when using matching strategies, as well as with plain encoding and decoding of hidden states. We observed that at lower layers, matching yields lower explained variance compared to encoding and decoding the previous hidden state. This can be explained by errors introduced by the matching algorithm at these layers.
@@ 6 Limitations and Conclusion
We introduced SAE Match, a novel data-free method for aligning Sparse Autoencoder (SAE) features across neural network layers, tackling the challenge of understanding feature evolution amid polysemanticity and feature superposition. By minimizing the mean squared error between folded SAE parameters—which integrate activation thresholds into the weights—we accounted for feature scale differences and improved matching accuracy. Experiments on the Gemma 2 language model validated our approach, showing effective feature matching across layers and that certain features persist over multiple layers. Additionally, we demonstrated that our method enables approximating hidden states across layers. This advances mechanistic interpretability by providing a new tool for analyzing feature dynamics and deepening our understanding of neural network behavior. We also showed that this approach can be applied not only to SAEs trained with the Gemma 2 model (see Figures 13, 19.
In the current version of the paper, we utilized only a scheme where we matched features based on all weights of the SAE (encoder, decoder, and biases). Future studies may include a more precise understanding of the behavior of different weights within the module. As we pointed out in Section 5.3, selecting appropriate SAEs based on their $l_{0}$ value is also a topic worth discussing. This value can be seen not only as a trade-off between sparsity and reconstruction quality, but also as a factor that affects matching features across layers. While in this work we focused on bijective matching to perform a step-by-step understanding of SAE features, it is evident that some features are not matched in this way. Therefore, one may want to explore more sophisticated methods to understand feature dynamics.
@@ Appendix A On the Matching with Different Sparsity
Figure 15: Left: The $l_{0}$ norms of the SAEs used in Neuronpedia and the most similar SAEs in GemmaScope (Lieberum et al., 2024). Right: Explained variance of the permutations. To analyze the behavior of the matching in the early layers, we conducted an additional experiment where we used a different SAE compared to Neuronpedia. We hypothesize that if two SAEs have different $l_{0}$ norms, then our approximation of the dynamics through permutations is less effective. To test whether our method performs better under these conditions, we selected SAEs with the most similar $l_{0}$ norms (see the red line in the left figure) and obtained permutations for this set. As shown in the right figure, reducing the "bumps" in $l_{0}$ between layers leads to a better explained variance metric. See Section 5.3 for more details.
Figure 16: Left: $\Delta L$ and Right: Explained Variance for various MSE quantiles.
Figure 17: Explained Variance dynamic for the different quntiles for Left 3rd layer Center 10th layer and Right 20th layer.
Appendix B MSE Quantiles and SAE Match as Layer Pruning
Quantile matching is performed via the following algorithm: let us denote $f_{q}^{-}$ as the activations of the features that have MSE lower than quantile $q$, and $f_{q}^{+}$ as the activations of the features with MSE higher than quantile $q$.
$\hat{h}=W_{dec}^{(t-1)}f_{q}^{+}+W_{dec}^{(t)}f_{q}^{-}+b_{dec}^{(t)}$ (6)
If quantile $q=1.0$, then $f_{q}^{+}=\overline{0}$ and therefore we get the original SAE Match. In contrast, if $q=0.0$, then $f_{q}^{-}=\overline{0}$ and we get the $D_{(t-1)}(E_{(t-1)}(h_{(t-1)}))$ modification with the Decoder’s bias equal to $b_{\text{dec}}^{(t)}$ instead of $b_{\text{dec}}^{(t-1)}$.
See Figures 16, 17 for the results. Generally, we observed that the lower the MSE value is, the larger the explained variance we could obtain, though for large MSE values we observed small variance in performance. From these results, we conclude that MSE reflects the semantic similarity of matched features.
Appendix C Evaluation Details
@@
0 width_16k/average_l0_102 13 width_16k/average_l0_83
1 width_16k/average_l0_108 14 width_16k/average_l0_83
@@
12 width_16k/average_l0_82 25 width_16k/average_l0_116
Table 1: Full list of SAEs we used for our experiments. See https://huggingface.co/google/gemma-scope-2b-pt-res for more details.
Appendix D Path Examples
@@
Figure 18: Examples of paths first, second, third and fourth features from 20th layer to 25th layer for Gemma-2-2B model.
Figure 19: Examples of paths first, second, third and fourth features from 10th layer to 28th layer for LLAMA-3.1-8B model.